OpenStudy (anonymous):

the length of a rectangle is 3 m less than five times the width. the area of the rectangle is 36 m^2. find the dimensions

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

lets assume the width to be W meters now try to write the length in terms of width with the information given

jhonyy9 (jhonyy9):

so write l-3=5w and what you know area of a rectangle is l*w l-3=5w l=5w+3 A=l*w A=(5w+3)w 36=5w^2 +3w 5w^2 +3w -36=0 solve it for w hope this will help you

OpenStudy (anonymous):

i got 3 for width